Among the 2,598,960 possible 5-card poker hands from a standard 52-card deck, how many contain the following cards?
a)Cards of more than one suit(complement of "all the same suit")

Answer by stanbon(75887)

Among the 2,598,960 possible 5-card poker hands from a standard 52-card deck, how many contain the following cards?
a)Cards of more than one suit(complement of "all the same suit")
-----
P(all 5 are the same suit) = (4*13C5)/52C5 = 0.001981
---
P(there are cards of more than one suit) = 1-(answer above)
================
Cheers,
Stan H.